OpenStudy (anonymous):

I need help factoring _7(x-2)_+_5_=_-5___ x-3 x x(x-3)

OpenStudy (anonymous):

its under each other

OpenStudy (anonymous):

like 7(x-2) x-3

OpenStudy (saifoo.khan):

\[\frac{7(x-2)}{x-3} + \frac5x = \frac{5}{x(x-3)}\]

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

\[\frac{(7x-14)x+5(x-3)}{x(x-3)}=\frac{5}{x(x-3)}\]

OpenStudy (saifoo.khan):

both the denomiators cancel out,

OpenStudy (saifoo.khan):

\[{7(x-2)x} + 5(x-3) = {5}\]

OpenStudy (saifoo.khan):

Solve further.

OpenStudy (anonymous):

thanks